A furniture store offers free setup and delivery services to all points within a 35 mi radius of its warehouse distribution center. If you love 19 mi east and 20 mi south of the warehouse, will you incur a delivery charge?


1
Expert's answer
2021-09-16T17:49:45-0400





Here A is the position of warehouse

C is the position of my house

AB = 19mi

BC = 20mi

AD = radius of circle = 35mi

Now by Pythagoras theorem we get the distance AC as follows.

AC = "\sqrt{AB² + BC²}"

= "\sqrt{19²+20²}"

= "\sqrt{361+400}"

= "\sqrt{761}"

= 27.59

So distance of my house from warehouse is approximately 27.59 mi which is less than the radius 35mi.

So no delivery charge will be incurred.